# Towards Stabilized and Efficient Diffusion Transformers through Long-Skip-Connections with Spectral Constraints

### v4

@@ Abstract
Diffusion Transformers (DiT) have emerged as a powerful architecture for image and video generation, offering superior quality and scalability. However, their practical application suffers from inherent dynamic feature instability, leading to error amplification during cached inference. Through systematic analysis, we identify the absence of long-range feature preservation mechanisms as the root cause of unstable feature propagation and perturbation sensitivity. To this end, we propose Skip-DiT, an image and video generative DiT variant enhanced with Long-Skip-Connections (LSCs) - the key efficiency component in U-Nets. Theoretical spectral norm and visualization analysis demonstrate how LSCs stabilize feature dynamics. Skip-DiT architecture and its stabilized dynamic feature enable an efficient statical caching mechanism that reuses deep features across timesteps while updating shallow components. Extensive experiments across the image and video generation tasks demonstrate that Skip-DiT achieves: (1) $\times4$ training acceleration and faster convergence, (2) 1.52×1.5-2$\times$1.5 - 2 × inference acceleration with negligible quality loss and high fidelity to the original output , outperforming existing DiT caching methods across various quantitative metrics. Our findings establish Long-Skip-Connections as critical architectural components for stable and efficient diffusion transformers. Codes are provided in the URL [^1] https://github.com/OpenSparseLLMs/Skip-DiT.
@@ 1 Introduction
In this study, we systematically analyze the feature distribution in DiT: ❶ We conduct a preliminary experiment that incorporates Long-Skip-Connections (LSCs) with DiT, and visualize the dynamic features of DiT with and without LSCs by adding perturbations in parameters and model inputs, the feature similarity landscape and plots validate the instability of vanilla DiT (w/o LSCs), which we formally characterize as Dynamic Feature Instability, indicating an uncontrolled spectral norm that affects both statistical stability, model robustness, and convergence rate. ❷ We then theoretically prove the superior stability and robustness of DiT with LSCs over vanilla DiT via spectral analysis.
@@
  • We introduce Skip-DiT, an enhanced DiT architecture incorporating Long-Skip-Connections between shallow and deep transformer blocks. Our approach provides theoretical guarantees for feature stability through spectral norm analysis and enables efficient feature caching at strategically selected timesteps.
  • We conduct extensive evaluations across different DiT base models, and various image and video generation benchmarks. Results of substantial training and inference speedup and nearly undamaged generation quality consistently endorse the effectiveness of Skip-DiT. For example, compared to vanilla DiT, Skip-DiT  achieves up to $\times4$ training acceleration with even better generation quality. When compared to state-of-the-art DiT caching techniques, Skip-DiT  delivers $1.5$-2×2$\times$2 × additional speedup while maintaining superior output fidelity.
@@ 2 Related Works
While Long Skip Connections (LSCs) have been incorporated into Diffusion Transformers (DiTs) by previous works like HunyuanDiT [14] and U-DiT [37], the underlying benefits of these connections have been largely unexamined. This paragraph clarifies the distinct advantages of our LSC implementation. Hunyuan DiT[14] is the first text-to-image model with LSCs, but a systematic exploration of their functional role has been absent from the literature. U-DiT [37] reshapes the Transformer into a U-Net structure [27] for efficiency, using token and feature downsampling for complexity reduction. In contrast, the primary motivation for Skip-DiT  is to improve feature stability and training convergence by using LSCs to optimize the spectral norm. This core design difference leads to a novel statistic-cached inference capability not present in other models. Furthermore, our feature fusion mechanism differs fundamentally. While U-DiT concatenates features, thereby doubling their dimension, Skip-DiT  employs a dimension-preserving fusion strategy that mixes recomputed and cached features from deep-to-shallow layers, enhancing robustness. Crucially, Skip-DiT  offers superior model compatibility. Its architecture, built upon native DiT blocks, allows for the direct adaptation of pretrained models. This is a significant advantage over U-DiT, which requires a full retraining process. For instance, we demonstrate this flexibility by successfully converting a pretrained DiT into our text-to-video model through continued training.
@@ 3.1 Preliminaries
Diffusion Transformer
In our study, we consider two types of DiT models processing different visual information: image DiT and video DiT. ❶ Image DiT model follows Vision Transformer, each block contains self-attention, cross-attention, and feed-forward networks (FFN), where the cross-attention module incorporates text and timestep conditions. ❷ Video DiT adopts dual-subblock architecture: the spatial Transformer subblock processes tokens within the same time frame, while the temporal subblock manages cross-frame relationships. These blocks alternate in sequence, with cross-attention integrating conditional inputs. A complete video DiT block pairs one spatial and one temporal component in an interleaved pattern, following [18, 53].
3.2 Dynamic Feature Instability in DiT
Visualizing Feature Stability of DiT
Let $\theta\in\mathbb{R}^{d}$ denote the original parameters of model $f(x;\theta)$. Given random perturbation vectors $\delta,\eta\in\mathbb{R}^{d}$, $\|\delta\|=\|\eta\|=\epsilon\|\theta\|$, with $10<\epsilon<$ 1 which is normalization coefficient. we construct perturbed parameters: $\theta^{\prime}=\theta+\alpha\delta+\beta\eta$, where $\alpha,\beta\in\mathbb{R}$ control perturbation magnitudes. The feature stability landscape for 3D visualization is defined as:
@@ 3.3 Integerating Long-Skip-Connection into DiT
To detail the DiT with Long-Skip-Connections(LSCs), we connect the shallow blocks to deep blocks with LSCs in DiT, named as Skip-DiT. Let $\bm{x}$ denote the input noise embedding, and $\bm{x}^{\prime}_{l}$ represents the output at the $l$-th layer of Skip-DiT. The architecture consists of $L$ sequential DiT blocks with LSCs. Each DiT block $\mathcal{F}_{\text{DiT}}^{l}$ at block $l$ processes the features as $\bm{x}^{\prime}=\mathcal{F}_{\text{DiT}}^{l}(\bm{x})$. The $i$-th skip branch ($i\in\{1\ldots L//2\}$) connects $i$-th block to ($L+1-i$)-th block, which can be denoted as $\mathcal{F}_{\text{Skip}}^{i}(\cdot,\cdot)$. Given output $\bm{x}^{\prime}_{i}$ from the start of the skip branch and $\bm{x}^{\prime}_{l}$ from the previous layer, the skip branch aggregates them to the input to $l$-th block as:
@@ 3.4 Spectral Analysis of Skip-DiT
Proof.  Define layer transformations for $L/2<l\leq L$ and their Jacobian matrices, where $\mathcal{T}$ denotes the Transformer block and $10<\alpha<$ 1 denotes normlized concatetion:
@@
This spectral norm reduction attenuates perturbation growth through $\|\Delta h_{L}\|\leq(\sigma_{\max}(J))^{L}\|\Delta h_{0}\|$, stabilizes gradient flow via $\|\nabla_{\theta}\mathcal{L}\|\leq\sigma_{\max}(J)\|\nabla_{h_{L}}\mathcal{L}\|$, and improves the Lipschitz constant $\mathrm{Lip}(M_{\mathrm{skip}})<\mathrm{Lip}(M_{\mathrm{vanilla}})$ with tight upper bound. This theoretical analysis aligns with our findings in Section 3.2, where Skip-DiT  are proven to be more feature stable and training efficient through comprehensive visualizations and preliminary experiments.
@@ 3.5 Cached Inference Mechanism
. As visualized in Figure 3, Skip-DiT  exhibits significantly higher feature similarity and smaller variance among examples compared to vanilla DiT, which employs FORA [30] and FasterDiffusion designed for DiT models. This demonstrates superior stability for cached inference. We attribute this improvement to the inherent feature stabilization and enhanced robustness. The latter can be theoretically grounded on $M$ defined in sec.3.4:
@@
4:$\bm{x^{\prime}}^{t}\leftarrow\mathcal{F}^{L-1}\circ\cdots\circ\mathcal{F}^{1}(\bm{x}^{t})$
@@
6:$\bm{x}^{t-1}=\mathcal{F}^{L}(\bm{x}^{t})$
7:for $k=1$ to $N-1$ do
8:     Local Timestep $t-k$ (Cached Inference):
@@
13:     if $t-k-1\in\mathcal{P}\leavevmode\nobreak\ ||\leavevmode\nobreak\ \epsilon>\text{threshold}$ then
@@ 4 Experiments
Table 1: Class-to-video generation performance. The definition of the cache step $n$ follows that in Table 2. For the UCF101 task, the scheduler is 50-steps-DDIM, while for the remaining tasks is 250-steps-DDPM. Latency and speedup are calculated on one A100 GPU. The notation $n=i$ represents cache interval $N$ in Algorithm 1. We highlight Baseline DiT models in grey, and the best metrics in bold.
@@
Latte 155.22 22.97 28.88 5.36 49.46 11.51 166.84 11.57 278.63 9.90 1.00$\times$
$\Delta$-DiT 161.62 25.33 25.80 4.46 51.70 11.67 188.39 12.09 226.10 8.09 1.22$\times$
FORA 160.52 23.52 27.23 4.64 52.90 11.96 198.56 13.68 240.26 9.00 1.10$\times$
PAB23 213.50 30.96 58.15 5.94 96.97 16.38 274.90 16.05 233.87 7.63 1.30$\times$
PAB35 1176.57 93.30 863.18 128.34 573.72 55.66 828.40 42.96 222.90 7.14 1.39$\times$
@@
original 141.30 23.78 20.62 4.32 49.21 11.92 163.03 13.55 290.05 10.02 1.00$\times$
$n=2$ 141.42 21.46 23.55 4.49 51.13 12.66 167.54 13.89 (0.34$\uparrow$) 180.68 6.40 1.56$\times$
$n=3$ 137.98 19.93 26.76 4.75 54.17 13.11 179.43 14.53 145.87 5.24 1.91$\times$
$n=4$ 143.00 19.03 30.19 5.18 57.36 13.77 188.44 14.38 125.99 4.57 2.19$\times$
$n=5$ 145.39 18.72 35.52 5.86 62.92 14.18 209.38 15.20 121.02 4.35 2.30$\times$
$n=6$ 151.77 18.78 42.41 6.42 68.96 15.16 208.04 15.78 111.07 4.12 2.43$\times$
@@ 4.1 Implementation Details
To start with, we add Long-Skip-Connections (LSCs) in DiT-XL [22] on the class-to-image task to prove the training efficiency and feature stability of  Skip-DiT. We employ Latte [18] as our base model in video generation and Hunyuan-DiT for text-to-image generation to further demonstrate the remarkable effectiveness of Skip-DiT  and its caching efficiency. Latte consists of spatial and temporal transformer blocks, suitable for both class-to-video and text-to-video tasks. Hunyuan-DiT [14], the first text-to-image DiT model with skip connections, is designed for text-to-image tasks. All DiT models are integrated with LSCs to evaluate the performance in both video and image generation tasks following the guidance in Section.3.3 and explore techniques for integrating LSCs into a pre-trained DiT model.
@@
For the class-to-image task, Skip-DiT is trained on 256$\times$256 ImageNet [7], following the dataset used by DiT-XL. In the class-to-video task, Skip-DiT  is trained on four datasets: FaceForensics [28], SkyTimelapse [42], UCF101 [34], and Taichi-HD [31]. Following the experimental settings in Latte, we extract 16-frame video clips from these datasets and resize all frames to a resolution of 256$\times$256. For the text-to-video task, the original Latte is trained on Webvid-10M [1] and Vimeo-2M [39], comprising approximately 330k text-video pairs in total. Considering that the resolution of Webvid-10M is lower than 512$\times$512, we only sample 330k text-video pairs from Vimeo-2M for training, with a resolution of 512$\times$512 and 16 frames.
@@ 4.2 Results on Video Generation Tasks
Table 2: Text-to-video generation performance. Inference latency and flops are evaluated with 1 A100 GPU. The notation $n=i$ represents the cache interval $N$ in Algorithm 1. Baseline DiT models are highlighted in grey, and the best metrics are emphasized in bold.
@@
Latte original 76.14 1587.25 27.11 1.00$\times$
T-GATE 75.68   ($\downarrow$0.46) 22.78 0.19 0.78 1470.72 24.15 1.12$\times$
FORA 76.06   ($\downarrow$0.08) 22.93 0.14 0.79 1341.72 24.21 1.19$\times$
$\Delta$-DiT 76.06   ($\downarrow$0.08) 24.01 0.17 0.81 1274.36 21.40 1.27$\times$
PAB235 73.77   ($\downarrow$2.37) 19.18 0.27 0.66 1288.08 23.24 1.24$\times$
PAB347 71.11   ($\downarrow$5.03) 18.20 0.32 0.63 1239.35 22.23 1.29$\times$
PAB469 70.20   ($\downarrow$5.94) 17.40 0.35 0.60 1210.11 21.60 1.33$\times$
AdaCache 76.07   ($\downarrow$0.07) 22.23 0.19 0.78 934.85 16.95 1.60$\times$
TeaCache 76.03   ($\downarrow$0.11) 23.05 0.17 0.80 862.52 15.75 1.72$\times$
Skip-DiT original 75.60 1648.13 28.72 1.00$\times$
@@
$n=2$ 75.51  ($\downarrow$0.09) 29.52 0.06 0.89 1127.83 19.28 1.49$\times$
$n=3$ 75.26  ($\downarrow$0.34) 27.46 0.09 0.85 974.80 16.67 1.72$\times$
$n=4$ 74.73  ($\downarrow$0.87) 25.97 0.13 0.81 882.98 15.12 1.90$\times$
@@
$n=2$ 75.36  ($\downarrow$0.24) 26.02 0.10 0.84 1066.62 18.25 1.57$\times$
$n=3$ 75.07  ($\downarrow$0.53) 22.85 0.18 0.76 852.38 14.88 1.93$\times$
$n=4$ 74.43  ($\downarrow$1.17) 22.08 0.22 0.73 760.56 13.03 2.20$\times$
@@
We compare the quantitative performance of Latte and Skip-DiT on four class-to-video tasks, as shown in Table 1. Skip-DiT consistently outperforms Latte in terms of FVD scores across all tasks while achieving comparable performance in FID scores, demonstrating its strong video generation capabilities. Furthermore, we observe that Skip-DiT  with cached inference significantly outperforms other caching methods across most metrics, incurring only an average loss of $2.37$ in the FVD score and $0.27$ in the FID score while achieving a 1.56$\times$ speedup.
@@ 4.3 Results on Image Generation Tasks
Table 3: Text-to-image generation performance. Latency and speedup are calculated on one A100 GPU. The notation $n=i$ represents cache interval $N$ in Algorithm 1. We highlight baseline DiT models (without caching) in grey, and the best metrics in bold.
@@
Table 4: Class-to-image generation performance. Speedups are calculated on an H100 GPU with a sample batch size of 8. The definition of $n=i$ is the same as other tables. The classifier-free guidance scale is set to 1.5. Skip-DiTis trained with 2.9M steps. We highlight baseline models in grey and the best metrics in bold.
Methods FID $\downarrow$ sFID $\downarrow$ IS $\uparrow$ PrecisionRecallSpeedup
@@
original 2.30 4.71 276.26 82.68 57.65 1.00$\times$
FORA 2.45 5.44 265.94 81.21 58.36 1.57$\times$
Delta-DiT 2.47 5.61 265.33 81.05 58.83 1.45$\times$
@@
Table 3 comprehensively compares Hunyuan-DiT and various caching methods. Notably, our caching mechanism achieves a 2.09$\times$ speedup without any degradation in FID or CLIP scores. Furthermore, it outperforms all other caching methods in terms of PSNR, LPIPS, and SSIM scores, consistently maintaining the highest performance even with a 1.96$\times$ speedup. These findings underscore the robustness and adaptability of our caching mechanism to image generation tasks. Additionally, the unusual improvement in FID and CLIP scores is due to: ❶ In HunyuanDiT’s standard evaluation pipeline, FID fluctuates due to zero-shot evaluation and resolution scaling ($1024\rightarrow 256$) prior to evaluation. ❷ The CLIP score is fluctuating within a reasonable range (-0.16, 0.2). These fluctuations are also observed in $\Delta$-DiT. The results on the class-to-image task shown in Table 4, Skip-DiT  also achieve the least loss while maintaining higher speedup.

### v5

@@ Abstract
Diffusion Transformers (DiT) have emerged as a powerful architecture for image and video generation, offering superior quality and scalability. However, their practical application suffers from inherent dynamic feature instability, leading to error amplification during cached inference. Through systematic analysis, we identify the absence of long-range feature preservation mechanisms as the root cause of unstable feature propagation and perturbation sensitivity. To this end, we propose Skip-DiT, an image and video generative DiT variant enhanced with Long-Skip-Connections (LSCs) - the key efficiency component in U-Nets. Theoretical spectral norm and visualization analysis demonstrate how LSCs stabilize feature dynamics. Skip-DiT architecture and its stabilized dynamic feature enable an efficient statical caching mechanism that reuses deep features across timesteps while updating shallow components. Extensive experiments across the image and video generation tasks demonstrate that Skip-DiT achieves: (1) $4.4\times$ training acceleration and faster convergence, (2) $1.5-2\times$ inference acceleration with negligible quality loss and high fidelity to the original output , outperforming existing DiT caching methods across various quantitative metrics. Our findings establish Long-Skip-Connections as critical architectural components for stable and efficient diffusion transformers. Codes are provided in the URL [^1] https://github.com/OpenSparseLLMs/Skip-DiT.
@@ 1 Introduction
In this study, we systematically analyze the feature distribution in DiT: ❶ We conduct a preliminary experiment that incorporates Long-Skip-Connections (LSCs) with DiT, and visualize the dynamic features of DiT with and without LSCs by adding perturbations in parameters and model inputs, the feature similarity landscape and plots validate the instability of vanilla DiT (w/o LSCs), which we formally characterize as Dynamic Feature Instability, indicating an uncontrolled spectral norm that affects both statistical stability, model robustness, and convergence rate. ❷ We then theoretically prove the superior stability and robustness of DiT with LSCs over vanilla DiT via spectral analysis
@@
  • We introduce Skip-DiT, an enhanced DiT architecture incorporating LSCs between shallow and deep transformer blocks. We provide theoretical guarantees for feature stability through spectral norm analysis and enable efficient feature caching at strategically selected timesteps.
  • Results of substantial training and inference speedup and nearly undamaged generation quality consistently endorse the effectiveness of Skip-DiT. For example, compared to vanilla DiT, Skip-DiT achieves up to $4.4\times$ training acceleration with even better generation quality. When compared to state-of-the-art DiT caching techniques, Skip-DiT delivers $1.5$-$2\times$ additional speedup while maintaining superior output fidelity.
@@ 2 Related Works
Although previous works like HunyuanDiT [14] and U-DiT [37] have incorporated Long Skip Connections (LSCs) into DiTs, a systematic analysis of their functional role has been absent. Our work provides this analysis, highlighting the distinct advantages of our implementation. We contrast our approach with U-DiT [37], which employs token and feature downsampling primarily for complexity reduction. Instead, our method uses deep-to-shallow LSCs to enhance the model’s spectral properties, boosting feature stability and training convergence. These connections preserve original dimensions, enabling an effective statistic-cached inference by mixing recomputed and cached features. A key advantage is that our method retains the native DiT block structure, permitting conversion from pre-trained models through continued training, unlike U-DiT which necessitates retraining from scratch.
@@ 3.1 Preliminaries
3.2 Dynamic Feature Instability in DiT
Visualizing Feature Stability of DiT
Let $\theta\in\mathbb{R}^{d}$ denote the original parameters of model $f(x;\theta)$. Given random perturbation vectors $\delta,\eta\in\mathbb{R}^{d}$, $\|\delta\|=\|\eta\|=\epsilon\|\theta\|$, with $0<\epsilon<1$ which is normalization coefficient. we construct perturbed parameters: $\theta^{\prime}=\theta+\alpha\delta+\beta\eta$, where $\alpha,\beta\in\mathbb{R}$ control perturbation magnitudes. The feature stability landscape for 3D visualization is defined as:
@@ 3.3 Integerating Long-Skip-Connection into DiT
To detail the DiT with Long-Skip-Connections(LSCs), we connect the shallow blocks to deep blocks with LSCs in DiT, named as Skip-DiT. Let $\bm{x}$ denote the input noise embedding, and $\bm{x}^{\prime}_{l}$ represents the output at the $l$-th layer of Skip-DiT. The architecture consists of $L$ sequential DiT blocks with LSCs. Each DiT block $\mathcal{F}_{\text{DiT}}^{l}$ at block $l$ processes the features as $\bm{x}^{\prime}=\mathcal{F}_{\text{DiT}}^{l}(\bm{x})$. The $i$-th skip branch ($i\in\{1\ldots L//2\}$) connects $i$-th block to (L+1iL+1-i)-th block, which can be denoted as $\mathcal{F}_{\text{Skip}}^{i}(\cdot,\cdot)$. Given output $\bm{x}^{\prime}_{i}$ from the start of the skip branch and $\bm{x}^{\prime}_{l}$ from the previous layer, the skip branch aggregates them to the input to $l$-th block as:
@@ 3.4 Spectral Analysis of Skip-DiT
Proof. Define layer transformations for $L/2<l\leq L$ and their Jacobian matrices, where $\mathcal{T}$ denotes the Transformer block and $0<\alpha<1$ denotes normlized concatetion:
@@
This spectral norm reduction attenuates perturbation growth through $\|\Delta h_{L}\|\leq(\sigma_{\max}(J))^{L}\|\Delta h_{0}\|$, stabilizes gradient flow via $\|\nabla_{\theta}\mathcal{L}\|\leq\sigma_{\max}(J)\|\nabla_{h_{L}}\mathcal{L}\|$. This theoretical analysis aligns with our findings in Section 3.2, where Skip-DiT is proven to be more feature stable and training efficient through visualizations and preliminary experiments.
@@ 3.5 Cached Inference Mechanism
As visualized in Figure 3, Skip-DiT exhibits significantly higher feature similarity and smaller variance among examples compared to vanilla DiT, which employs FORA [30] and FasterDiffusion designed for DiT models. This demonstrates superior stability for cached inference. We attribute this to the inherent feature stabilization and enhanced robustness. The latter can be theoretically grounded on $M$ defined in Section 3.4:
@@
[…]
@@
[…]
7:for $k=1$ to N1N-1 do
8:  Local Timestep tkt-k (Cached Inference):
@@
13:  if $t-k-1\in\mathcal{P}\penalty 10000\ ||\penalty 10000\ \epsilon>\text{threshold}$ then
@@ 4 Experiments
Table 1: Class-to-video generation performance. The definition of the cache step $n$ follows that in Table 2. For the UCF101 task, the scheduler is 50-steps-DDIM, while for the remaining tasks is 250-steps-DDPM. Latency and speedup are calculated on one A100 GPU. The notation $n=i$ represents cache interval NN in Algorithm 1. We highlight Baseline DiT models in grey, and the best metrics in bold.
@@
[…]
@@
[…]
@@ 4.1 Implementation Details
To start with, we add LSCs in DiT-XL [23] on the class-to-image task to prove the training efficiency and feature stability of Skip-DiT. Then we employ Latte [17] as our base model in video generation and Hunyuan-DiT for text-to-image generation to further demonstrate the remarkable effectiveness of Skip-DiT and its caching efficiency. All DiT models are integrated with LSCs to evaluate the performance in both video and image generation tasks following the guidance in Section.3.3 and explore techniques for integrating LSCs into a pre-trained DiT model.
@@
For the class-to-image task, Skip-DiT is trained on $\times256$ ImageNet [7]. In the class-to-video task, Skip-DiT is trained on four datasets: FaceForensics [28], SkyTimelapse [42], UCF101 [34], and Taichi-HD [31]. Following the experimental settings in Latte, we extract 16-frame video clips from these datasets and resize all frames to a resolution of $\times256$. For the text-to-video task, the original Latte is trained on Webvid [1] and Vimeo [39], comprising approximately 330k text-video pairs in total. Considering that the resolution of Webvid is lower than $\times512$, we only sample 330k text-video pairs from Vimeo for training, with a resolution of $\times512$ and 16 frames.
@@ 4.2 Results on Video Generation Tasks
Table 2: Text-to-video generation performance. Inference latency and flops are evaluated with 1 A100 GPU. The notation $n=i$ represents the cache interval NN in Algorithm 1. Baseline DiT models are highlighted in grey, and the best metrics are emphasized in bold.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4.3 Results on Image Generation Tasks
Table 3: Text-to-image generation performance. Latency and speedup are calculated on one A100 GPU. The notation $n=i$ represents cache interval NN in Algorithm 1. We highlight baseline DiT models (without caching) in grey, and the best metrics in bold.
@@
Table 4: Class-to-image generation performance. Speedups are calculated on an H100 GPU with a sample batch size of 8.
Methods FID $\downarrow$ sFID $\downarrow$ IS $\uparrow$ Precision% Recall% Speedup
@@
[…]
@@
[…]